OpenStudy (anonymous):

Which does not show a direct variation between x and y? A. y = 0.5x B. y = 9/x C. y = x/9 D. y = 2x

OpenStudy (phi):

direct variation means y = k x where k is some number. remember that x/9 can be written as \[ \frac{1}{9} x \] which means k= 1/9 so you have direct variation

OpenStudy (phi):

all the direct variations will be number times x which choice is not "number times x" ?

OpenStudy (anonymous):

which one?

OpenStudy (phi):

is A number times x ?

OpenStudy (anonymous):

idk but is it a b c or d

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

do you know which one it is?

OpenStudy (phi):

are you saying you can't pick out the choices that have a number times x ?

OpenStudy (phi):

for choice A, is 0.5 a number ? is x an "x" ? is 0.5 x a number times x ?

OpenStudy (phi):

what about choice D ? is 2 a number ? is 2 x a number times x ?

OpenStudy (phi):

in choice C, is 1/9 a number ? is 1/9 x a number times x ?

OpenStudy (anonymous):

is it choice c?

OpenStudy (anonymous):

? @phi

OpenStudy (anonymous):

@phi hello is it c?

OpenStudy (phi):

in choice C, is 1/9 a number ? is 1/9 x a number times x ?

OpenStudy (phi):

is 1/9 a number ? YES 1/9 x means 1/9 times x that is direct variation

OpenStudy (anonymous):

so its c? yes or no

OpenStudy (phi):

First is 1/9 a number ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

direct variation is a number times x 1/9 * x or 1/9 x or x/9 all mean the same thing.... a number 1/9 times x so it is direct variation you want the choice that is not direct variation

OpenStudy (anonymous):

ok thanks so its c right yes or no this is like my 3rd or 4th time!

OpenStudy (phi):

re-read what I posted.

OpenStudy (anonymous):

please is it c or not

OpenStudy (anonymous):

it is !!!!!!!!!!!!!!!

OpenStudy (phi):

you have to read more carefully. I wrote multiple times that choice C IS DIRECT variation you want the choice that is NOT direct variation if choice C is DIRECT variation, it cannot be the answer.

OpenStudy (anonymous):

..........................

OpenStudy (phi):

you could find the answer by eliminating all the choices that are direct variation A y = 0.5 x <-- direct B.y = 9/x or 9 (1/x) C.y = x/9 or y= (1/9) x<-- direct D. y = 2 x<-- direct

OpenStudy (anonymous):

its b..............